Ed weighs 10 kilograms more than his sister Debbie. Their combined weight is 120 kilograms. How many kilograms does Ed weigh?
Answer by JulietG(1812)

E + D = 120 (Their combined weight is 120 kilograms.)
E = D + 10 (Ed weighs 10 kilograms more than his sister Debbie)
Substitute the known value of E from the second equation into the first.
(D+10) + D = 120
2D + 10 = 120
Subtract 10 from each side
2D = 110
Divide each side by 2
D = 55
Debbie weighs 55kg and Ed weighs 65kg